OpenStudy (e.cociuba):

For each quadratic equation, find the: a.) Roots b.) y-intercept c.) Vertex y= 4(x-4)^2 +16

OpenStudy (e.cociuba):

Ok so for the Roots I got (2,6) is that correct? Also, I got 80 for the y-intercept. Is that also correct?

OpenStudy (e.cociuba):

I'm having a bit hard time finding out the vertex, so I also need help on that.

OpenStudy (jhannybean):

for y-intercept plug in x=0. \[y= 4(0-4)^2 +16 = 4(-4)^2+16 = 80\] the y-intercept is correct

OpenStudy (e.cociuba):

Yea I thought so to :) Thanks!

OpenStudy (jhannybean):

to find the roots you solve for your x-values, \[4(x-4)^2=16\]\[(x-4)^2 = 4\]\[\sqrt{(x-4)^2}= \sqrt{4}\]\[x-4 =2\] \[x=6\]

OpenStudy (jhannybean):

plug x back into the equation to find y

OpenStudy (e.cociuba):

Yup so you get (2,6)

OpenStudy (e.cociuba):

I'm just having a hard time finding the vertex.

OpenStudy (jhannybean):

oh wait, i did something wrong :( i think you do this: \[(x-4)^2 = 4\] \[x-4 =-2\] and \[x-4=2\] so yes you get 2 and 6 and tehn for the vertex, let me see...

OpenStudy (e.cociuba):

Yup. Ok I just need help figuring the vertex. I don't just want the answer, I wan't to know how to do it too, u no?

OpenStudy (jhannybean):

Mmhmm. First you need to find the Axis of symmetry, the line that cuts the parabola in half. And the point on the graph that is represented by the x and y-value. In order to find it,you have the formula \[x= \frac{ -b }{ 2a }\] so you have to expand the formula \[y=4(x-4)^2+16\]

OpenStudy (jhannybean):

So expand it and tell me what you get :)

OpenStudy (e.cociuba):

Ok so I usually do this in my calculator but it's messed up so I'm gunna do it your way :) Ok so were using the Quadratic Formula right?

OpenStudy (jhannybean):

\[y=4(x^2-8x+16)+16 = 4x^2-32x+64+16 =4x^2-32x+80\]

OpenStudy (e.cociuba):

It's a bit confusing...

OpenStudy (jhannybean):

So now what you have to do is apply \[x=\frac{ -b }{ 2a }\] with your "b" and "a" values from this equation we just expanded, and find your x-value. Once you find your x-value you plug that x-value back into\[y=4(x-4)^2+16\] to get your y-value The (x,) value you will derive will represent the coordinate on which the axis of symmetry your parabola lies on, which is your vertex. :)

OpenStudy (e.cociuba):

I'm sorry but I'm seriously not getting it. It's ugghh a bit frustrating.

OpenStudy (jhannybean):

what are you not getting? Which portion precisely.

OpenStudy (e.cociuba):

The whole x = -b/2a It looks like you didn't apply it to your equation^^^

OpenStudy (jhannybean):

What do you mean didn't apply it to my equation? i didn't write out my steps. I can write out my step if you like. Do you have the answer to what the vertex is?

OpenStudy (e.cociuba):

Ok so I usually do this in my calc but its not responding. I never learned how to do this by hand yet so that's y I'm lost. :3

OpenStudy (jhannybean):

Oh, okay. Well just like you use the quadratic to find your values of x, (usually you get 2 vaules using a quadratic), you can use the equation \[x=\frac{ -b }{ 2a }\] to find THE X-VALUE of the vertex coordinate. once you find this value, you plug it back into your original (or expanded equation) to get THE Y-VALUE of the coordinate of the vertex

OpenStudy (jhannybean):

Also, there is this video if you don't understand my explanation :) http://www.virtualnerd.com/algebra-1/quadratic-equations-functions/graphing/graph-basics/vertex-example There, the quadratic expression had already been expanded, which i advised you to do.

OpenStudy (e.cociuba):

Ok thanks for all your patience and work! I appreciate the help! :)

OpenStudy (jhannybean):

Does it make sense at all? :(

jimthompson5910 (jim_thompson5910):

why expand? the equation y= 4(x-4)^2 +16 is already in vertex form

OpenStudy (e.cociuba):

Not in the least. Sorry. But I'm gunna check that video out and ask my professor for one on one time. Thanks though!

jimthompson5910 (jim_thompson5910):

there's a reason why it's called that y= a(x-h)^2 +k has the vertex (h,k)

jimthompson5910 (jim_thompson5910):

ex: y= 2(x-7)^2 +33 has the vertex (7,33)

OpenStudy (jhannybean):

Ohh... I was working AROUND that because i forgot about the (h,k)...

OpenStudy (e.cociuba):

So the (4,16) is the vertex?

OpenStudy (jhannybean):

Yes.

OpenStudy (e.cociuba):

Oh that's it?

OpenStudy (jhannybean):

Thanks Jim :)

jimthompson5910 (jim_thompson5910):

np

OpenStudy (jhannybean):

That's it! But i tried explaining HOW to get the vertex as well... sorry or confusing you!

OpenStudy (e.cociuba):

Well I didn't no it was that simple. My professor always makes things harder than they have to be. :)

OpenStudy (jhannybean):

I summon Jim when I cannot get a point across. He's a lifesaver.

OpenStudy (e.cociuba):

Haha np girl. Your extremely smart by the way! Thanks for the help! :)

OpenStudy (jhannybean):

Oh thanks ^_^ and np!

OpenStudy (e.cociuba):

:)

OpenStudy (e.cociuba):

Thanks @jim_thompson5910 also! I appreciate the help guys! :)

jimthompson5910 (jim_thompson5910):

yw